OpenStudy (anonymous):

f(x)= 1/2 abs value x-2 abs value

OpenStudy (helder_edwin):

\[ \large f(x)=\frac{1}{2}|x-2| \] ??

OpenStudy (anonymous):

yes i need the domain and range

OpenStudy (helder_edwin):

well there no roots (of even power), no division, and no logarithms. so the domain is \(\mathbb{R}\). now since the absolute value is non-negative then \[ \large |x-2|\geq0 \] \[ \large f(x)=\frac{1}{2}|x-2|\geq\frac{1}{2}0=0 \] so the range is \([0,+\infty)\)

OpenStudy (anonymous):

thank you!

OpenStudy (helder_edwin):

u r welcome